OpenStudy (anonymous):

Another Limit question. 1/x+1 - 1/2 On x-1 Lim x -> 1

OpenStudy (anonymous):

\[\frac{ 1/x+1 - 1/2 }{ x-1 }\]

OpenStudy (anonymous):

(1/x+1) - (1/2)

OpenStudy (anonymous):

i am going to guess it is \[\frac{\frac{1}{x-1}-\frac{1}{2}}{x-1}\]

OpenStudy (anonymous):

yes, a little hard to format haha

OpenStudy (anonymous):

nope \[\huge \frac{\frac{1}{x+1}-\frac{1}{2}}{x-1}\]

OpenStudy (anonymous):

... mb

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

what i wrote i mean

OpenStudy (anonymous):

Yes now it is

OpenStudy (anonymous):

you want to replace x by 1 but you cannot , because you will get \(\frac{0}{0}\) that means your goal is to cancel the factor of \(x-1\) top and bottom ] to do this you need to simplify this compound fraction q

OpenStudy (anonymous):

it is algebra all the way first off do the subtraction in the numerator subtract \[\frac{1}{x+1}-\frac{1}{2}\]

OpenStudy (anonymous):

So you multiply the denominator to match them up? (2)(X+1)

OpenStudy (anonymous):

yes, the denominator will be \(2(x+1)\)

OpenStudy (anonymous):

leave it in factored form is fine

OpenStudy (anonymous):

\[\frac{ x+1 }{ 2(x+1) } - \frac{ 2 }{ 2(x+1) }\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

x + 1 cancel out and 2 remains lim x -> 1 = 2

OpenStudy (anonymous):

ending up the the \(x-1\) up top that you want to cancel with the bottom

OpenStudy (anonymous):

crap yes

OpenStudy (anonymous):

whoa hold the phone you still have to subract

OpenStudy (anonymous):

you still have a compound fraction as well

OpenStudy (anonymous):

\[\frac{ x+1 }{ 2(x+1) } - \frac{ 2 }{ 2(x+1) }=\frac{x-1}{2(x+1)}\] and that is just then numerator

OpenStudy (anonymous):

but now it is easy divide by \(x-1\) and you are left only with \[\frac{1}{2(x+1)}\] and now you are free to replace \(x\) by \(1\)

OpenStudy (anonymous):

Oh okay, I was wondering how you got x - 1 for a sec. But thanks! This really helps :)

OpenStudy (anonymous):

yw